Summary of the 3mm GMVA Session May, 7-12, 2009 --------------------------------------------------- In this session all stations participated, with the exception of Metsahovi, which is waiting for a new 86 GHz receiver. An ftp fringe check done on for PV, PdB and Eff on a 20-second data segment (00:36:00 to 00:36:20, 3C273) on May 8 showed good results between PdB and PV. Missing tracks detected for Effelsberg were traced back to a Mk5 wiring problem, which was immediately fixed after it became obvious. The whole session was effected by bad weather, with rain, heavy clouds in the second half of the session in EB and high wind (> 20 m/s) on Pico Veleta. In EB the weather was good on the first 2.5 days, but degraded afterwards. Pico Veleta had good conditions only on the first 1.5 days, and was weathered out due to too high wind from May 9 until May 12, 14UT. PdB was much less affected by bad weather and observed almost continuously (except for sun-avoidance and some period of not so good weather at the end). On May 11, 22 UT PdB had to stop observing due to high wind and on May 12 observed only a few of scans. From the VLBA logs, one can expect good data for all 8 stations (approx. 9 % downtime). NL and BR were affected by some rain. Some more details: ------------------- Effelsberg: no major problem. Stable pointing, improved by U. Bach (two more new pointing constants) May 7 day 127 clear sky May 8 day 128 thunderstorm 0-6 UT, wind and rain Mk5 cabling problem fixed 12:45 UT May 9 day 129 good weather until noon, later increasing cloud coverage May 10 day 130 good weather in the night and morning, later partly cloudy, rain starts at 22 UT May 11 day 131 rain 2-18h overcast after 18h00, but can see some sources May 12 day 132 overcast, then rain, better after 18h00 Pico Veleta: observed 7 May 100% of scheduled 2h. (22:00 to 24:00) 8 May 50% 12 h (00 to 4:30 and 16 to 24) 9 May 8% 3.5 h. (00 to 2:30 and 4:45 to 5:45) 10 May 0% high wind 11 May 0% high wind 12 May 40% 10h. 14:00 to 24:00 Total observed 28h of 122 (23%) total lost time about 77% due to wind. Plateau de Bure: May 7 ok, Tsys 110K, phases correct, wind low no losses May 8 ok, Tsys about 110K, some light clouds, low wind 15h30: rain, stopped observing 16 UT: resume, Tsys around 135K. Foggy passages and clouds lost 9 scans 13:30 UT - 15:40 UT (rain/snow) May 9 observing all night under good conditions, Tsys is about 100K 11h40: stop due to high wind, rain showers 14h15: resume, but sun avoidance for 3C111 and 0420-014 lost 2 scans 07:45 UT - 07:55 UT (program crash) lost 4 scans 10:33 UT - 11:07 UT (high wind) lost 7 scans 11:45 UT - 14:10 UT (high wind) May 10 high clouds, wind close to limits, observing until morning sun avoidance during day time 19h00 UT: resume observing on 3C273 lost 1 scan 06:30 UT - 06:40 UT (system stuck) lost 2 scans 10:15 UT - 10:40 UT (high wind) May 11 good conditions 22h10 UT : the wind has been steadily rising over the last hours, wind stop lost 2 scans 03:15 UT - 03:40 UT (rain/snow) lost 7 scans 22:15 UT - 23:59 UT (high wind) May 12 until 14 UT: only ~10 scans observed, high wind day time: sun avoidance 18h30-22:25 UT: observed OJ287, then wind stop lost 10 scans 00:00 UT - 02:25 UT (high wind) lost 9 scans 04:15 UT - 06:25 UT (high wind) lost 26 scans 07:15 UT - 13:55 UT (high wind) lost 2 scans 18:00 UT - 18:25 UT (high wind) lost 12 scans 20:30 UT - 23:25 UT (high wind) Total: scheduled: 598 scans lost: 93 scans, mainly due to weather sun avoid.: 43 scans observed: 462 scans (83% of possible scans) Onsala: observed, no detailed information yet, looks ok day 127 ok Tsys 260 day 128 rain 6-13 UT, 22 UT rain again day 129 ok, changing weather 250-600 K day 130 ok, Tsys 240-400 K day 131 ok, moderate Tsys ~300-500 K day 132 ok, changing weather, better after 9h, bad again after 22 UT Metsahovi: no receiver, no data VLBA: Weather from daily webcam inspection May 8 13 UT: NL overcast, rains, other stations ok May 9 0 UT: NL overcast, other stations clear and ok 10 UT: BR clouds, NL light clouds 16 UT: FD cloudy, others good May 10 16 UT: MK clouds, other good May 11 16 UT: all ok 19 UT: BR clouds May 12 12 UT: all ok 25 UT: NL overcast, PT clouds, others ok VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C091A Observer: Krichbaum, T. Project type: OTHER Obs filename: c091acrd.* Date/Day: 2009MAY07/127 Ants Scheduled: NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Paul Padilla GPS delta, UT1-UTC, and VClock are all ok 2357 Begin 2357-2352 %FD 3mm cryo cycling, up to 35K at times. CRYO 2357-2352 %LA 3mm Tsys values running about twice the 24844 50 3597 expected value. 2357-2122 *KP pointing offsets found to have been 24780 100 1285 affected by azimuth wheel work last week. AZ encoder offset and collimation offsets affected. 0000 Date is 2009MAY08/128. 0011-0013 *PT HC2 FIFO Formatter Error, Cleared FMTR 100 2 0022 %OV windy 0025-0026 *PT HC2 FIFO Formatter Error, Cleared FMTR 100 1 0112-0113 *PT HC2 FIFO Formatter Error, Cleared FMTR 100 1 0138 %NL Raining 0259-0300 *PT HC2 FIFO Formatter error, cleared itself FMTR 100 1 0338-0339 *PT HC2 FIFO Formatter error FMTR 100 1 0452-0454 *PT HC2 FIFO Formatter error FMTR 100 2 0525-0526 *PT HC2 FIFO Formatter error FMTR 100 1 0535 New Operator is Alan Kerr 0537-0540 *MK focus drive fault. FRM 100 3 0609-0614 *PT HC2 FIFO error. FMTR 100 5 0713-0715 *PT HC2 FIFO error. FMTR 100 2 0741-0744 *PT HC2 FIFO error. FMTR 100 3 0750-0800 *MK focus drive fault. would not clear. FRM 100 10 0755-0758 *PT HC2 FIFO error. FMTR 100 3 0844-0850 *MK focus drive fault. FRM 100 6 0852-0858 *PT HC2 FIFO error. FMTR 100 6 0927-0936 *NL precipitation. 0955-0958 *PT HC2 FIFO error. FMTR 100 3 1007-1012 *PT HC2 FIFO error. FMTR 100 5 1107-1237 %NL short power failure. GPS error, unclearable 24832 100 90 FIFO errors (FIFO x empty). Site Tech called out. 1109-1118 *NL precipitation. 1115-1123 *MK focus drive fault. wouldn't clear. FRM 100 8 1123-1127 *PT HC2 FIFO error. FMTR 100 4 1420 Operator is A. Sowinski. 1426-1438 *NL Site tech fixing UPS, requires computers 24832 100 12 (VME and Mark5) to be shut down. 1450 KP wea: clear, dry. 1518 OV wea: clear, dry. 1447-1449 *PT HC2 FIFO errors. FMTR 100 2 1703-1704 *PT Power glitch. COM 100 1 1734-1738 *NL Site tech finishing fixing UPS. 24832 100 4 1800-1858 *NL Power cable to D-rack failed. 24835 100 58 2122 *Rebooting KP VME to make sure new pointing 24780 equation parameters are properly loaded. During a setup scan so no data lost. 2145 New Operator: Paul Padilla 2145-0555 %PT occasional Formatter errors, mostly on FMTR non-recording scans 2247-2249 *NL autostowed, power glitch COM 100 2 0000 Date is 2009MAY09/129. 0255-0320 %HN raining 0550 New Operator is Alan Kerr 0555 PT&KP occasional FIFO errors. If cleared in under 2 minutes, it will not be logged. 1028-1031 *MK focus drive fault. FRM 100 3 1350 New operator is Heidi Medlin 1530-1534 KP comm down, still observing 1825-2123 FD BUTIL MCB error. Station building sensors 24838 reading ****. Not affecting data. 0000 Date is 2009MAY10/130. 0137 New operator: T. Perreault 1345 New operator is Heidi Medlin 1418-1424 *PT autostowed, power glitch COM 100 6 1903-1906 *KP autostowed, power glitch COM 100 3 2324-2326 *KP autostowed, power glitch COM 100 2 2336 *FD site tech reports isolated T-storms in area 0000 Date is 2009MAY11/131. 0135 New operator: T. Perreault 0922 OV wea: 35+ mph wind gust 1252 PT wea: clear, no precipitation, dry 1344 New operator is Heidi Medlin 1423-1809 MK put into standby while source is below horizon need to repoint at 18:16 today. 1424-1438 *FD put into standby for site tech OTHER 100 14 1522-1525 *PT autostowed, power glitch COM 100 3 1631-1632 *KP mark5 comm lost TAPE 100 1 2157 New operator is Betty Ragan. 2300 OV wea: cloudy, dry. 0000 Date is 2009MAY12/132. 0546 New Operator: Paul Padilla 0910-0913 *PT HC2 Formatter error FMTR 100 3 1058-1100 *KP autostowed, power glitch COM 100 2 1137 %OV windy 1342 New operator is Heidi Medlin 1610-1625 *KP stowed for site tech to look at frm FRM 100 15 1637-1639 *BR rebooted itself, possibly due to a power glitch COM 100 2 1944-2200 MK put into stand by while source is below horizon will need to repoint at 22:00. 2132-2133 *PT autostowed, commercial power glitch. Repointed. COM 100 1 2156-2159 *PT HC2 formatter error. FMTR 100 3 2243 New operator is Betty Ragan. 2244 OV wea: clear, dry, moderate but gusty winds. 2352 End. Downtime Summary: Total downtime : 5176 min Percentage downtime of observing: 9.0% Average downtime per hour : 5.4 min Total scheduled observing time (# Antennas): 57560 min (8) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard.